A builder lifts eight slabs from the ground on to the back of a lorry 1.5 m high.
The total time taken is 48 s and each slab weighs 200 N.
How much useful power does the builder produce?
1
Expert's answer
2019-04-29T09:42:41-0400

The power by definition

"P=\frac{\rm{work\:done}}{\rm{time}}=\frac{mgh}{t}"

"=\frac{8\times 200\:\rm{N}\times 1.5\:\rm{m}}{48\:\rm{s}}=50\:\rm{W}"